a chemist has two solution of sulfuri acid a 20% solution and an 80% solution. how much of each should be used ro obtain 100 liters of a 62% solution?
Answer by macston(5194)

.
Let x=liters of 20% solution
0.20(x)+0.80(100L-x)=.62(100L)
0.20x+80L-0.80x=62L
-0.60x+80L=62L Add 0.60 x to each side.
80L=0.60x+62L Subtract 62L from each side.
18L=0.60x Divide each side by 0.60
30L=x
ANSWER 1: He should use 30 liters of 20% solution.
100L-30L=70L
ANSWER 2: He should use 70 liters of 80% solution.
.
CHECK:
.
0.20(30)+0.80(70)=0.62(100)
6L+56L=62L
62L=62L